OpenStudy (anonymous):

Consider the function f(x) = 2 x^3 - 6 x^2 - 48 x + 9 on the interval [ -6 , 8 ]. Find the average or mean slope of the function on this intervalBy the Mean Value Theorem, we know there exists a c in the open interval ( -6 , 8 ) such that f'( c) is equal to this mean slope. For this problem, there are two values of c that work.

OpenStudy (anonymous):

The mean slope is 44 i already figured it out and it is correct.

OpenStudy (anonymous):

I need to find both values of c

OpenStudy (anonymous):

You know how to do derivatives?

OpenStudy (anonymous):

yes should be 6x^2 - 12x -48 right?

OpenStudy (anonymous):

Yeup now factor this

OpenStudy (anonymous):

set it equal to zero?

OpenStudy (anonymous):

Yes. What do you get by setting it to zero.

OpenStudy (anonymous):

6(x-4)(x+2) = 0

OpenStudy (anonymous):

And your values of c are...

OpenStudy (anonymous):

4 and -2?

OpenStudy (anonymous):

both those are not the correct answers

OpenStudy (anonymous):

Nope. You're not done yet.

OpenStudy (anonymous):

plug in for y?

OpenStudy (anonymous):

you mean for x. I think you have to do that in the original equation. I'm not sure :/

OpenStudy (anonymous):

wouldn't i set the derivative equal to the average slope?

OpenStudy (anonymous):

so f'x = 44

OpenStudy (anonymous):

because its asking for 2 points c with the same slope i think

OpenStudy (anonymous):

I think there is one value of c that equals to mean slope.

OpenStudy (anonymous):

so 6(x-4)(x+2) = 44?

OpenStudy (anonymous):

i get x= 5 1/3, 11 1/3 but those do not work either

OpenStudy (anonymous):

Hmm I'm not sure anymore. I thought this was gonna work. I'll work on this.

OpenStudy (agent0smith):

so 6(x-4)(x+2) = 44? i get x= 5 1/3, 11 1/3 Those aren't correct. Expand it all out, simplify, use quadratic formula.

OpenStudy (anonymous):

C is 9, it is given by the function I think.

OpenStudy (anonymous):

I think agent is right. Use the quadratic formula.

OpenStudy (anonymous):

quadratic formula worked! -3.04, and 5.04

OpenStudy (anonymous):

awesome!!

OpenStudy (anonymous):

wish i could best response both of you >.<

OpenStudy (agent0smith):

I just checked your equation on wolfram alpha, and both solutions had square roots.

OpenStudy (anonymous):

So maybe the negative one isn't right or do you think she did the quadratic equation wrong?

OpenStudy (anonymous):

Her answers are right. Just exact.

OpenStudy (agent0smith):

btw by "I just checked your equation on wolfram alpha, and both solutions had square roots." i meant i checked these two: x= 5 1/3, 11 1/3 the other two were right.

OpenStudy (anonymous):

Yeah